OpenStudy (anonymous):

Challenge! Ready? the length of the hypotenuse of a 30 degrees-60 degrees-90 degrees triangle is 18. What is the perimeter?

OpenStudy (anonymous):

This should be easy.

OpenStudy (wolf1728):

In a 30 60 90 triangle, short side = hypotenuse / 2 and the medium side = hypotenuse *sq root(3)/2 short side =9 and medium side = 0.8660254038 * 18 = 15.5884572681 So add all 3 for perimeter (info about 30 60 90 triangles here: http://www.1728.org/trig2.htm

OpenStudy (anonymous):

Hint, this is going to be at the end.

OpenStudy (anonymous):



OpenStudy (wolf1728):

You need the perimeter right?

OpenStudy (anonymous):

Yes. You guys are doing it correctly.

OpenStudy (anonymous):

This what it will look like. : __+__

OpenStudy (hwyl):

+_+

OpenStudy (anonymous):



OpenStudy (anonymous):

What numbers go in the blanks?

OpenStudy (wolf1728):

Okay we know the sides are 18 + 9 + 18*sq root(3) /2 or 18 + 9 + 9*sq root(3)

OpenStudy (anonymous):

Ikr, it's fun.

OpenStudy (anonymous):

Seriously who cares if I give them out? I can do whatever as a user. Like as if that picture is actually of you; you are prob. a guy.

OpenStudy (anonymous):

So you and your fake pic. Look at me using the same judgmental logic.

OpenStudy (wolf1728):

Lollygirl - you are here to get answers too aren't you? :-)

OpenStudy (anonymous):

For the last time, I am as innocent as a flower. Can I not have fun?

OpenStudy (wolf1728):

Lollygirl do I get a medal?

OpenStudy (hwyl):

challenge question, are you ready? 2+2 = 5

OpenStudy (anonymous):

Lol.

OpenStudy (wolf1728):

thanks :-)

OpenStudy (anonymous):

You are doing really well, wolf, it is know wonder you have 94 for problem solving.